A spring with a spring constant of 5 N/m is compressed by a force of 1.2 Newtons a distance of 0.11m. What is the total elastic potential energy stored in the compressed spring?



1
Expert's answer
2021-06-12T10:42:30-0400
"PE_s=\dfrac{1}{2}kx^2,""PE_s=\dfrac{1}{2}\cdot5\ \dfrac{N}{m}\cdot(0.11\ m)^2=0.03\ J."